OpenStudy (anonymous):

solve 4tan^2x+5tanx=6

OpenStudy (campbell_st):

well this is a quadratic equation \[4\tan^2(x) + 5(x) - 6 = 0\]

OpenStudy (mathstudent55):

Bring the 6 over to the left side. Then solve the quadratic equation for tan x. Then solve the trig part.

OpenStudy (campbell_st):

it can be factored.... to (4tan(x) + 8)(4tan(x) - 3) ----------------------- = 0 4 remove a common factor from the 1st binomial 4(tan(x) + 2)(4tan(x) - 3) ---------------------- = 0 4 cancel the common factor then solve for x

OpenStudy (anonymous):

wait can you reexplain it to me please... I don't understand how you factored it

OpenStudy (campbell_st):

ok... well a simple method for factoring non-monic quadratic is multiply a and c \[ax^2 + bx + c=0\] so in your question it was 4 x - 6 = -24 now find the factors that add to b, or in your question 5.......8 and - 3 then you write (ax + factor 1)(ax + factor 2) ------------------------- =0 a and you'll find that there are common factors in the binomial that cancel with the denominator... it works everytime

OpenStudy (anonymous):

oh okay...thank you so much